OpenStudy (anonymous):

which is equal to log25 5?

OpenStudy (anonymous):

log5/log25? log25/log5?

OpenStudy (anonymous):

or log1/5

OpenStudy (anonymous):

log 25 5?!

OpenStudy (anonymous):

Is it?

OpenStudy (anonymous):

I know in mathematica 25 5 = 25*5

OpenStudy (anonymous):

log 25 5 = log25+log5 or is 25 the base ? \[\log _{25} 5\] = 1/2